OpenStudy (anonymous):

Can somebody explain how my solutions manual got 11.8673 for 60 divided by 30 squared?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

60/900 = 6/90 = 2/30

OpenStudy (anonymous):

by being retarded.

OpenStudy (anonymous):

.0666666666666................................ thats what you will get

OpenStudy (anonymous):

@joemath we call that "differently abled" now

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

by being differently abled.

OpenStudy (anonymous):

So my college algebra textbook screwed up on this one?

OpenStudy (anonymous):

Did not think such a thing was feasible, but I suppose it is..:P